OpenStudy (anonymous):

how would you solve the inequality - g(x)>1/2 when g(x)=sinx

OpenStudy (anonymous):

Solved, sin-1(1/2) = pi/6 pi-pi/6= 5/6 pi Therefore the answer = pi/6 <x< 5/6 pi